U was also ANSWER the following: what are the orders of the following functions F(P,9) - plog(P) + pq I F(P,9) - plog(P) + p2 can someone help me in writing BigOnotations for this ?
Answer #1

a)

Big-O notation of f(p, q)  = O(pq)

b)

Big-O notation of f(p, q)  = O((p + q)^2)

c)

Big-O notation of f(p, q)  = O(p log(q))